### v3

# jina-embeddings-v3: Multilingual Embeddings With Task LoRA

Abstract
We introduce jina-embeddings-v3, a novel text embedding model with 570 million parameters, achieves state-of-the-art performance on multilingual data and long-context retrieval tasks, supporting context lengths of up to 8192 tokens. The model includes a set of task-specific Low-Rank Adaptation (LoRA) adapters to generate high-quality embeddings for query-document retrieval, clustering, classification, and text matching. Evaluation on the MTEB benchmark shows that jina-embeddings-v3 outperforms the latest proprietary embeddings from OpenAI and Cohere on English tasks, while achieving superior performance compared to multilingual-e5-large-instruct across all multilingual tasks. With a default output dimension of 1024, users can flexibly reduce the embedding dimensions to as low as 32 without compromising performance, enabled by Matryoshka Representation Learning.
**footnotetext: Equal contribution.
1 Introduction
Text embedding models represent documents as high-dimensional vectors, converting semantic relationships between documents into spatial relationships between vectors. These models are fundamental to neural information retrieval and have been widely adopted across various domains of NLP and IR research and applications. Text embeddings are utilized in diverse downstream tasks such as classification, retrieval, and clustering. Notably, they have gained significant traction in building Retrieval-Augmented Generation (RAG) systems, where they serve as the primary technique in the retrieval step.
A major limitation of traditional embedding models is that, despite being named as general-purpose, they often require fine-tuning for specific tasks Jiao et al. (2020) and frequently struggle with common failure cases Gao et al. (2021). To address this, recent research has increasingly focused on leveraging large language models (LLMs) as the backbone for general-purpose embedding generation, capitalizing on their ability to efficiently handle multiple languages and tasks Jiang et al. (2024). However, with model sizes typically reaching 7 billion parameters, deploying these models in real-world applications poses significant challenges. Furthermore, the marginal improvements in evaluation metrics offered by LLM-based embeddings, compared to encoder-only embedding models, render them a less practical choice for many use cases.
This paper introduces jina-embeddings-v3, a novel text embedding model with 570 million parameters, optimized for multilingual data, long-context retrieval, and high performance across multiple tasks. Evaluation on the MTEB benchmark demonstrates that jina-embeddings-v3 not only significantly improves upon its predecessor, jina-embeddings-v2 Günther et al. (2023) and its bilingual variants Mohr et al. (2024), but also outperforms the latest proprietary embeddings from OpenAI and Cohere on English tasks, while surpassing multilingual-e5-large-instruct across all multilingual tasks. Additionally, compared to LLM-based embeddings such as e5-mistral-7b-instruct, which has a parameter size of 7.1 billion (12x larger) and an output dimension of 4096 (4x larger) but offers only a 1% improvement on MTEB English tasks, jina-embeddings-v3 is a far more cost-efficient solution, making it more suitable for production and on-edge computing. The key contributions of this paper are:
  • Task-specific optimization with LoRA: We demonstrate that LoRA adapters Hu et al. (2021) effectively generate task-specific embeddings, outperforming prior instruction-based approaches.
  • Patching retrieval failures with synthetic data: A qualitative analysis identified four common types of retrieval failures. We mitigated these issues by incorporating synthetic training data, thereby improving model robustness on edge cases.
  • Integration of latest techniques: Our model incorporates several key advancements, including Matryoshka Representation Learning Kusupati et al. (2022), instruction tuning Wei et al. (2022); Su et al. (2023), and long-context retrieval Günther et al. (2023).
Section 2 provides an overview of prior research relevant to the objectives of this paper. Section 3 presents the architecture of jina-embeddings-v3 in detail. The training procedure is described in Section 4. In Section 5, we conduct a thorough multilingual evaluation, including ablation studies that offer insights into the impact of our architectural and training decisions.
2 Related Work
2.1 General Text Embeddings
In recent years, significant progress has been made in the field of text embeddings, largely driven by the emergence of transformer-based pre-trained language models that capture the underlying semantics of language effectively Devlin et al. (2019). However, these models are predominantly trained with a masked language modeling (MLM) objective, which is not optimal for generating high-quality text embeddings. To overcome this limitation, recent approaches have focused on fine-tuning and extending these models specifically for embedding tasks Reimers and Gurevych (2019).
A key advancement in this area is the development of multi-stage and multi-task fine-tuning strategies that incorporate weakly-supervised contrastive training Wang et al. (2022); Günther et al. (2023); Mohr et al. (2024). These methods improve the versatility of embeddings, enabling models to perform well across a diverse range of applications and tasks, as opposed to models trained solely on semantic textual similarity datasets.
Furthermore, techniques such as AliBi Press et al. (2022) and RoPE Su et al. (2024) have enabled models like jina-embeddings-v2 Günther et al. (2023) to handle longer sequences, up to 8192 tokens, by replacing absolute positional encoding with relative encoding methods. To make embeddings more compact, Matryoshka Representational Learning (MRL) Kusupati et al. (2022) enables the truncation of embeddings without compromising performance on downstream tasks by modifying the loss function used during training.
2.2 Multilingual Embedding Models
One of the earliest multilingual transformer models is Multilingual BERT (mBERT) Devlin et al. (2019), trained on 104 languages. This was followed by XLM Conneau and Lample (2019) and XLM-RoBERTa (XLM-R) Conneau et al. (2020), which utilized parallel data during training. Wang et al. (2024) extends this work by fine-tuning XLM-R on high-quality multilingual labeled datasets and applying knowledge distillation from a cross-encoder to further improve embedding quality. Similarly, Chen et al. (2024a) introduced BGE M3, another XLM-R-based model that supports longer sequences. The authors extended XLM-R’s maximum sequence length to 8192 tokens, continued pre-training with the RetroMAE method Xiao et al. (2022), and fine-tuned it contrastively using a novel multi-CLS pooling strategy. mGTE Zhang et al. (2024) also builds on XLM-R, incorporating RoPE positional embeddings Su et al. (2024).
Another approach leverages LLMs for multilingual embeddings Zhang et al. (2023); Wang et al. (2023), benefiting from their extensive language support and diverse training data. However, LLMs are computationally inefficient due to their larger size, making them less practical for many applications. To address this, Lee et al. (2024a) generate and relabel training data to distill knowledge from LLMs into a compact encoder model, avoiding the need for direct fine-tuning of the larger LLMs.
2.3 Task-Specific Embedding Models
Previous research has highlighted limitations in training models to produce generic embedding vectors that perform well across various use cases and domains. For example, Wang et al. (2022) observed that in asymmetric retrieval tasks, such as question answering and typical information retrieval, models perform better by appending distinct prefixes to queries and documents before encoding. While the E5 models from this work employ a single prefix for all queries and another for all documents, Su et al. (2023) introduced more complex instructions to encode additional information about relevance in retrieval tasks and the domain of the data.
Hu et al. (2021) propose a technique that uses lightweight LoRA layers to fine-tune LLMs. By freezing the original model weights, this approach significantly improves training efficiency. More importantly, deploying multiple fine-tuned instances becomes feasible, as the LoRA typically require less than 1% of the memory needed for the original model weights. However, to the best of our knowledge, this technique has not yet been explored as an alternative to instruction-based methods in embedding training.
3 Model Architecture
Figure 1: The architecture of jina-embeddings-v3 is based on the XLM-RoBERTa model, with several key modifications. FlashAttention 2 is integrated for enhanced computational efficiency, while RoPE extends support for sequences up to 8192 tokens. Task-specific LoRA adapters are introduced to optimize embeddings for various tasks. The model’s input consists of two parts: the text, which is the long document to be embedded, and the task type. jina-embeddings-v3 supports four tasks and implements five adapters to choose from: retrieval.query and retrieval.passage for query and passage embeddings in asymmetric retrieval tasks, separation for clustering and reranking tasks, classification for classification tasks, and text-matching for tasks involving semantic similarity, such as STS or symmetric retrieval.
The architecture of jina-embeddings-v3 is depicted in Figure 1. To implement the backbone architecture, we adapt the XLM-RoBERTa model with modifications that (1) enable effective encoding of long text sequences, (2) allow task-specific encoding of embeddings, and (3) increase model efficiency. jina-embeddings-v3 retains the original XLM-RoBERTa tokenizer.
As outlined in Table 1, jina-embeddings-v3 is larger than jina-embeddings-v2, but significantly smaller than embedding models fine-tuned from LLMs Lee et al. (2024b); Wei et al. (2022). Importantly, the LoRA adapters account for less than 3% of the total parameters, adding minimal overhead. To further enhance performance and reduce memory consumption, we utilize FlashAttention 2 Dao , support activation checkpointing, and employ the DeepSpeed framework Rasley et al. (2020) for efficient distributed training.
Parameters Max input tokens Max output dim. Layers Vocabulary Attention Impl. Pooling
Base w/ adapters
559M 572M 8192 1024 24 250K FlashAttention2 Mean
Table 1: Model specification of jina-embeddings-v3. The maximum output dimension is 1024, but with Matryoshka Representation Learning (MRL), users can choose any dimension lower than 1024, such as 16 or 32, allowing for a trade-off between space efficiency and performance. This trade-off is studied in Table 7.
To handle long text sequences, we replace absolute positional embeddings with Rotary Position Embeddings (RoPE) Su et al. (2024), which use a rotation matrix to encode absolute positions while embedding relative positional dependencies directly within the self-attention mechanism. We also experimented with extending positional encodings  [^1] https://spaces.ac.cn/archives/7947 (Accessed 08-27-2024) as done in the BGE M3 model Chen et al. (2024b), but observed poor performance on tasks involving long texts. This could be attributed to differences in training data and pooling strategies, as we trained primarily on short texts and used mean pooling instead of multi-CLS pooling.
Xiong et al. (2024) demonstrated that increasing the base frequency parameter of rotary positional embeddings enhances performance on long-text tasks, while Zhang et al. (2024) adjusted the rotary base frequency during training on short sequences to generalize better on longer sequences. We found that setting a rotary base frequency of 10,000 during training and adjusting it to 20,000 during inference improves performance on long-text tasks without degrading performance on short-text tasks.
Embeddings can be used for various downstream tasks, including clustering, retrieval, and classification, each requiring different interpretations of the representation space resulting in different similarity metrics. Asymmetric retrieval tasks, for instance, benefit from encoding queries and documents differently. Wang et al. (2022) suggest that using distinct instructions for queries and documents improves the effectiveness of embedding models in such tasks.
However, writing effective task-specific instructions is non-trivial. As an alternative, we employ task-specific LoRA adapters. The embedding and linear layers within the multi-head attention mechanism are equipped with low-rank decomposition matrices of rank 4. These task-specific LoRA adapters are loaded alongside the model weights and are dynamically selected based on the input task type. Each text input in the batch is associated with a task descriptor, represented as an integer corresponding to the LoRA adapter’s ID.
4 Training Method
We initialize the model using the weights of the original XLM-RoBERTa model. However, the model’s original MLM objective is not fully aligned with our training objectives due to the changes in positional embedding methods. Despite this, we observe that initializing with pre-trained weights leads to faster convergence during pre-training compared to random initialization.
Our training paradigm consists of three stages, as is common for training text embedding models:
  1. I
    Pre-Training: We perform standard MLM training using large multilingual text corpora. The model is initialized with XLM-RoBERTa weights to expedite pre-training and avoid training from scratch.
  2. II
    Fine-Tuning for Embedding Tasks: To learn how to encode a text passage into a single vector representation, we follow the approach outlined in Günther et al. (2023). This method incorporates a pooling layer into the transformer model to aggregate token representations into a single embedding vector and fine-tunes the model on pairs of semantically related texts.
  3. III
    Training Task-Specific Adapters: We train five LoRA adapters for four different tasks using dedicated datasets and task-specific loss functions to optimize performance for each use case.
4.1 Pre-Training
After initialization, the model is trained using the MLM objective with whole word masking Devlin et al. (2019). At this stage, we only train the transformer model, excluding the LoRA adapters and pooling layer.
To support multilingual tasks, the training data is drawn from the CulturaX corpus Nguyen et al. (2023), which includes data from 89 languages, with English comprising approximately 20% of the dataset. During training, each batch contains data for only a single language, but we rotate the language between batches.
For long-context support, we first train for $100,000$ steps on text sequences that are truncated to 512 tokens, followed by an additional $60,000$ steps using on text sequences truncated to 8192 tokens and a reduced batch size. The details are provided in Appendix A1.
To enhance the model’s ability to represent long text documents, we train with a low rotary base value, increasing it during inference as described in Section 1. However, we observed that the model’s ability to encode long documents still lagged behind models such as jina-embeddings-v2. To address this, we extended the training on long-text data, which resulted in improved performance on long-text retrieval tasks such as NarrativeQA. See Section 5.3 for further details.
4.2 Fine-Tuning for the Embedding Task
After pre-training, we fine-tune the model to encode a text sequence into a single vector representation. Following the Sentence-BERT approach Reimers and Gurevych (2019), we augment the model with a mean pooling layer to aggregate the semantics from all output token vectors into a single vector representation. The fine-tuning procedure follows Mohr et al. (2024), where the model is trained on text pairs using a bi-directional InfoNCE (van den Oord et al., 2018) loss, $\mathcal{L}_{\mathrm{pairs}}$:
$\mathcal{L}_{\mathrm{pairs}}(B):=\mathcal{L}_{\mathrm{NCE}}(B)+\mathcal{L}_{\mathrm{NCE}}(B^{\dagger})$ (1)
defined on a batch $B=((p_{1},q_{1}),\ldots,(p_{k},q_{k}))$ of $k$ pairs, and $B^{\dagger}=((q_{1},p_{1}),\ldots,(q_{k},p_{k}))$ (obtained from $B$ by swapping the order of pairs). $\mathcal{L}_{\mathrm{NCE}}$ denotes the following loss function:
$\mathcal{L}_{\mathrm{NCE}}(B):=-\sum_{(x_{i},y_{i})\in B}\ln\frac{e^{s(x_{i},y_{i})/\tau}}{\sum\limits_{i^{\prime}=1}^{k}e^{s(x_{i},y_{i^{\prime}})/\tau}}$ (2)
The training data consists of over one billion text pairs, drawn from more than 300 distinct sub-datasets, each representing specific domains in various languages. During training, the data loader constructs each batch by sampling a specific sub-dataset, ensuring that only text pairs from that dataset and language are included.
For data preparation, we follow the same methodology as our previous work Mohr et al. (2024), with an additional filtering step. This filter removes pairs where at least 80% of the words (minimum of four) in the shorter text are sub-strings of the longer text. This filtering step increases the difficulty of the training and encourages the model to focus less on syntactic overlap.
As in the pre-training phase, we begin training on short text pairs, followed by further training on longer texts using a larger sequence length but a reduced batch size. In this phase, we use only a subset of the datasets containing sufficiently long texts.
4.3 Training Task-Specific Adapters
Related work on embedding training Wang et al. (2022); Xiao et al. (2023) introduces an additional generic training phase following the pair-wise training phase. This phase incorporates high-quality data from various task types to optimize model performance across a range of downstream use cases. In this stage, recent approaches use task-specific instructions to help the model distinguish between different tasks and domains, as discussed in Section 2.3.
However, this approach complicates the usage, as users must learn task-specific instructions (i.e., prompts) that align with the model’s behavior or "vibe." While this offers flexibility, it also makes the model’s behavior harder to predict. In contrast, we train five distinct LoRA adapters for four well-defined task types as defined in Table 2. These tasks are trained independently, with the base model’s weights kept frozen. For query-document retrieval, two adapters are trained jointly: one for queries and one for passages. During inference, users can select the appropriate adapter based on their downstream task and input role, ensuring optimal embeddings for their specific use case.
task value Task Description
retrieval.passage Embedding documents in a query-document retrieval task
retrieval.query Embedding queries in a query-document retrieval task
separation Clustering documents, visualizing a corpus
classification Text classification
text-matching Semantic text similarity, general symmetric retrieval, recommendation, finding similar items, deduplication
Table 2: Supported tasks of jina-embeddings-v3, each corresponding to a LoRA adapter and trained independently, except for retrieval.passage and retrieval.query, which are trained jointly.
4.3.1 Classification Adapter
The classification adapter generates embeddings that are effective for training downstream classification models, particularly logistic regression classifiers. To train the adapter, we employ the classification training method proposed for the Gecko embedding model Lee et al. (2024a). Specifically, we select datasets covering a range of common classification tasks, including sentiment analysis, intent classification, and article categorization.
From each dataset, we construct tuples consisting of two text values from the same class $(q,p)$ and seven text values from different classes $(n_{1},\dots,n_{7})$, resulting in a tuple of nine text values $(q,p,n_{1},\dots,n_{7})$. The model is trained to assign a high cosine similarity to the embeddings of $q$ and $p$, while enforcing low cosine similarity between $q$ and all $n_{i}$. Each batch is composed of tuples sampled from a single dataset.
We employ an extended version of the InfoNCE loss $\mathcal{L}_{\text{triplet}}$ described in our previous work Günther et al. (2023) to take into account these additional negative samples.
$\mathcal{L}_{\text{triplet}}(B):=$
$\;\;\;\;\;\mathbb{E}_{r\sim B}\Bigg{[}-\ln\frac{e^{s(q,p)/\tau}}{\sum\limits_{i=1}^{k}\Big{[}e^{s(q,p_{i})/\tau}+\sum\limits_{j=1}^{m}e^{s(q,n_{j,i})/\tau}\Big{]}}\Bigg{]}$
$\,+\mathbb{E}_{r\sim B}\Bigg{[}-\ln\frac{e^{s(p,q)/\tau}}{\sum\limits_{i=1}^{k}e^{s(p,q_{i})/\tau}}\Bigg{]}$
$\text{with}\;r=(q,p,n_{1},\ldots,n_{m}).$ (3)
When using this loss function, texts from the same class as text $q_{i}$ that occur in the same batch as negatives for a different text $q_{j}$ $(i\neq j)$ are also treated as negatives. This introduces the problem of false negatives.
To address this, Lee et al. (2024a) propose appending a unique ID specific to each tuple $r$ to its corresponding text values. This allows the model to easily differentiate between text values from different tuples, enabling it to focus on separating text values within the same tuple in a batch.
4.3.2 Text Matching Adapter
This adapter is trained to produce embeddings that quantify the similarity between two text values. It is applicable for tasks such as semantic textual similarity (STS) and retrieval tasks where there is no clear distinction between query and target text values. An example of such a retrieval task is duplicate detection, where text values from a corpus are compared against each other. In these cases, the “query” and “corpus” texts are treated symmetrically.
To train this adapter, we use the CoSent loss function: $\mathcal{L}_{\mathrm{co}}$ [^2] https://github.com/bojone/CoSENT (Accessed: 08-28-2024), as previously employed by Li and Li (2024):
$\mathcal{L}_{\mathrm{co}}(B):=\ln\Big{[}1+\sum\limits_{\begin{subarray}{c}(q_{1},p_{1}),\\ (q_{2},p_{2})\in B\end{subarray}}\frac{e^{s(q_{2},p_{2})}-e^{s(q_{1},p_{1})}}{\tau}\Big{]}$
$\text{where}\;\zeta(q_{1},p_{1})>\zeta(q_{2},p_{2})$ (4)
The CoSent loss operates on two pairs of text values, $(q_{1},p_{1})$ and $(q_{2},p_{2})$, which are constructed from the batch by selecting combinations of four text values where the ground truth similarity $\zeta$ is provided in the training dataset, and $\zeta(q_{1},p_{1})$ is greater than $\zeta(q_{2},p_{2})$.
To train the model with this objective, we use data from semantic textual similarity (STS) training datasets such as STS12 Agirre et al. (2012) and SICK Marelli et al. (2014). These datasets consist of triplets $(q_{i},p_{i},t_{i})\in D$, where $(q_{i},p_{i})$ are text pairs and $t_{i}$ is the corresponding relevance score. A batch $B$ is constructed by selecting text values from a given number of triplets, with the ground truth similarity defined as $\zeta(q_{i},p_{i})=t_{i}$.
To enhance the model’s performance across languages, we translate the STS12 and SICK datasets into multiple languages using machine translation models, i.e. WMT19 Ng et al. (2019) and MADLAD-3B Kudugunta et al. (2023). While training on STS datasets is highly effective, obtaining large volumes of this data is challenging due to the human annotation required. As a result, we have incorporated various natural language inference (NLI) datasets into the training process.
During each training step, either an STS or NLI dataset is selected, and the batch is constructed solely from the chosen dataset, using the appropriate loss function. In other words, each batch contains text values from only one specific dataset.
For the relevant hyperparameters, see Appendix A1.
4.3.3 Asymmetric Retrieval Adapters
As discussed in Section 2.3, asymmetric retrieval tasks, such as question answering and traditional information retrieval, perform better with separate encoding mechanisms for queries and documents. In this work, we follow the method proposed by Wang et al. (2022), using two distinct prefixes, but further separate the encoding processes by employing two specialized adapters, which are trained jointly. A detailed ablation study demonstrating the effectiveness of this approach is presented in Section 5.5.2.
Similar to previous works Wang et al. (2022); Li et al. (2023); Günther et al. (2023), we use datasets containing hard negatives, such as MS-MARCO Bajaj et al. (2016) and Natural Questions (NQ) Kwiatkowski et al. (2019), to train the model to focus on subtle distinctions and to differentiate between relevant and similar but irrelevant documents. For retrieval training datasets without annotated negatives, we apply hard negative mining as outlined in (Ren et al., 2021; Wang et al., 2022), leveraging embedding models like BGE-large Xiao et al. (2023) and BM25 Robertson et al. (2009).
To incorporate the mined negatives into the training process, we apply the $\mathcal{L}_{\text{triplet}}$ loss function as seen in Equation (3).
4.3.4 Failure Analysis for Asymmetric Retrieval
Since our jina-embeddings-v2 models were trained on similar data to jina-embeddings-v3, we conducted a failure analysis to identify issues common to models trained on these datasets. From this analysis, we identified the following points affecting retrieval tasks:
  • F1.
    Misleading Syntactic Similarities: Documents with high syntactic similarity to the query are often favored over gold/relevant documents with lower syntactic overlap.
  • F2.
    Misinterpretation of Named Entities: Named entities are frequently not recognized as such, leading to documents being marked as relevant based on partial matches (e.g., "Sofia Albert" vs. "Albert Stone"). This occurs especially with proper nouns that have alternative, more common meanings (e.g., the novel title "The Company" vs. "the company").
  • F3.
    No Understanding of Polar Questions: Complex yes-no (polar) questions are not handled effectively. As a result, the model retrieves documents with related content that do not necessarily answer the query.
  • F4.
    Preference for Low-Quality Documents: jina-embeddings-v2 and many other embedding models do not account for document quality, focusing solely on similarity and relevance. Consequently, low-quality documents (short, repetitive, or uninformative) that mention query terms are often retrieved but do not provide satisfactory answers.
To mitigate F1–F3, we crafted prompts to generate synthetic text examples targeting these specific failure cases. Each example consists of a query text, a preferred answer, and seven negative examples modeling the failure case.
For F4, we leveraged two preference learning datasets: oasst1333https://huggingface.co/datasets/OpenAssistant/oasst1 (Accessed: 30-08-24) and oasst2444https://huggingface.co/datasets/OpenAssistant/oasst2 (Accessed: 30-08-24) from the Open Assistant project Köpf et al. (2024). These datasets contain questions and answers generated by LLMs with quality scores (0-1) based on human judgments.
We converted these datasets into hard negative training data by selecting queries with at least two answers. The highest-quality answer was treated as a positive match, while answers with at least a 0.3-point lower quality score were treated as negatives. If fewer than seven negatives were identified, additional negatives were randomly selected from other queries.
The effectiveness of training the retrieval adapter on this data is evaluated in Section 5.4.
4.3.5 Separation Adapter
The separation adapter is designed to perform well on clustering and reranking tasks. It is trained to distinguish between text values belonging to the same group and those from different groups. For reranking tasks, the adapter separates relevant from irrelevant documents based on a query’s information need. In clustering tasks, groups of texts are provided, and after calculating the embeddings and applying a clustering algorithm (e.g., k-means), the resulting clusters should reflect the correct groupings.
To train the adapter for this objective, we employ a variant of the CoSent loss $\mathcal{L}_{\mathrm{co}}$, introduced in Equation (4.3.2). The training data consists of batches $B^{\prime}$ made up of tuples $(x,l)\in B^{\prime}$, where $x$ is a text value and $l$ is its label. To form a batch of text pairs compatible with $\mathcal{L}_{\mathrm{co}}$, we generate all pairwise combinations of text values that share the same label $l_{i}$ in the batch. The separation loss is then defined as follows:
$\mathcal{L}_{\mathrm{sep}}(B^{\prime}):=\mathcal{L}_{\mathrm{co}}(B)$
$B=\{(x_{i},x_{j})|\exists l:(x_{i},l),(x_{j},l)\in B^{\prime}\}$ (5)
Since we have a limited amount of training data in this format, we observed that incorporating additional data from the pair training stage (Section 4.2) into the training mix improves performance. We follow the same schema as used for the text matching adapter (Section 4.3.2), where a specific dataset is sampled at each training step, and the corresponding loss function is applied.
Further details on the training data and hyperparameters are provided in Appendix A1.
5 Evaluation
In this section, we evaluate the performance of our model at various stages and conduct ablation studies on key architectural modifications. We begin by assessing the multilingual backbone model on a small subset of MTEB tasks in Section 5.1.
Next, we present a comprehensive evaluation of embedding tasks in Section 5.2, where our model is tested on a variety of MTEB tasks, both monolingual (English) and multilingual. Section 5.3 reports the model’s performance on the LongEmbed MTEB evaluation, followed by an analysis of previously identified retrieval failure cases in Section 5.4. Lastly, Section 5.5 presents the ablation studies conducted on MRL and the retrieval adapter.
Tasks Jina-XLM-R XLM-R mBERT
Multi-/Cross-lingual Tasks*
STS22STS 64.52 64.80 60.59
STS17STS 77.09 75.13 69.90
PawsXPC 59.75 59.93 56.65
Average 67.12 66.62 62.38
EN Tasks
QuoraRetrievalRT 84.88 83.49 77.81
STS12STS 71.29 70.26 70.44
STS13STS 81.55 80.89 75.69
STS14STS 73.29 71.90 69.94
STS15STS 83.17 81.03 77.99
STS16STS 79.54 77.96 76.27
STS17STS 84.48 84.00 80.35
STS22STS 65.98 66.93 64.83
TRECCOVIDRT 60.28 42.96 35.31
Average 76.05 73.38 69.18
*Results per task are averages across all languages
for that task.
Table 3: Evaluation of multilingual pre-trained models after short embedding training on pair-wise data.
Model Average CF CL PC RR RT STS SM
jina-embeddings-v2-base-en 60.38 73.45 41.73 85.38 56.98 47.87 80.70 31.60 English
jina-embeddings-v3 65.52 82.58 45.27 84.01 58.13 53.87 85.80 29.71
text-embedding-3-large 64.60# 75.45 49.01 - 59.16 55.44 - -
multilingual-e5-large-instruct 64.41 77.56 47.10 86.19 58.58 52.47 84.78 30.39
Cohere-embed-multilingual-v3.0 64.01 76.01 46.60 86.15 57.86 53.84 83.15 30.99
jina-embeddings-v2-base-(zh/es/de)* 60.54 65.69 39.36 82.95 66.57 58.24 66.60 - Multi-L
jina-embeddings-v3 64.44 71.46 46.71 76.91 63.98 57.98 69.83 -
multilingual-e5-large 59.58 65.22 42.12 76.95 63.40 52.37 64.65 -
multilingual-e5-large-instruct 64.25 67.45 52.12 77.79 69.02 58.38 68.77 -
CF: Classification Accuracy   CL: Clustering $\mathcal{V}$ measure  PC: Pair Classification Average Precision
RR: Reranking MAP  RT: Retrieval nDCG@10  STS: Sentence Similarity Spearman Correlation
SM: Summarization Spearman Correlation  (Scores in %)
*: jina-embeddings-v2 bilingual model suite, only evaluated on Chinese, Spanish, and German tasks.
#: The average MTEB score for text-embedding-3-large model as reported in the announcement blog post
Table 4: Performance of multilingual text embedding models on MTEB tasks as averages.
5.1 Performance of Jina-XLM-RoBERTa
We evaluate the Jina-XLM-RoBERTa model on a subset of English and multi-/cross-lingual MTEB tasks, conducting a comparative analysis against established multilingual models, specifically mBERT Devlin et al. (2019) and XLM-RoBERTa Conneau et al. (2020), which are widely used as backbones for multilingual embedding models.
For this experiment, we follow the same training setup described in Section 4.2, but limit training to 1000 steps on a single GPU node, processing approximately 2 million pairs. Adapter tuning is excluded from this phase. As shown in Table 3, our model outperforms both XLM-R and mBERT across all tasks, achieving an average of 76.05% on monolingual English tasks and 67.12% on multi-/cross-lingual tasks. These results reinforce our decision to continue training XLM-R for multilingual applications.
5.2 Performance on MTEB
Table 4 summarizes the performance of various multilingual text embedding models across different MTEB tasks, divided into English and multilingual sections. jina-embeddings-v3 performs competitively, particularly in monolingual English tasks, where it achieves the highest Classification Accuracy (CF) score of 82.58 and the top Sentence Similarity (STS) score of 85.80, demonstrating its robustness across both languages and tasks. Full evaluation results per task can be found in Appendix A3. When averaging across all tasks, jina-embeddings-v3 scores 65.52, outperforming models such as text-embedding-3-large, multilingual-e5-large-instruct, and Cohere-embed-multilingual-v3.0. This indicates that jina-embeddings-v3 maintains strong monolingual English performance while being trained on a wide variety of languages.
Consulting the English MTEB leaderboard, we plotted the performance of the top 100 embedding models against their parameter sizes in Figure 7. Notably, embedding models built on large language models (LLMs) perform only marginally better than jina-embeddings-v3 but entail significantly higher complexity and computational costs in real-world applications. For example, e5-mistral-7b-instruct achieves an average score of 66.63 across all 56 English MTEB tasks (approximately 1.03% higher than jina-embeddings-v3), but produces embeddings with a dimension of 4096 and has a parameter size of 7.1 billion. In contrast, jina-embeddings-v3 offers a more practical solution, with an embedding dimension of 1024 (or lower, using MRL with a manageable performance trade-off, as discussed in Section 5.5.1) and only 570 million parameters.
For multilingual performance, Table 4 presents weighted averages across a wide selection of multilingual and cross-lingual MTEB tasks. The specific tasks, along with an explanation of which adapter was used for each task, are detailed in Appendix A7, A6, A8, A5, A9, and A4. Note that jina-embeddings-v2 refers to our bilingual model suite (jina-embeddings-v2-base-zh, jina-embeddings-v2-base-es, jina-embeddings-v2-base-de), which are evaluated only on Chinese, Spanish, and German monolingual and cross-lingual tasks, excluding all other tasks. Thus, the average score for jina-embeddings-v2* on the pair-classification tasks reflects performance on Chinese tasks alone. A complete report on pair-classification can be found in Appendix A6. We do not report scores for text-embedding-3-large and Cohere-embed-multilingual-v3.0 as these models were not evaluated on the full range of multilingual and cross-lingual MTEB tasks. However, our model outperforms multilingual-e5-large in all multilingual tasks except reranking and approaches the performance of multilingual-e5-large-instruct.
For jina-embeddings-v3, all classification and pair-classification tasks are evaluated using the classification adapter, all STS tasks and three retrieval tasks (ArguAna, CQADupstackRetrieval, and QuoraRetrieval) are evaluated using the text matching adapter, all other retrieval tasks are evaluated using the retrieval adapter, and all clustering and reranking tasks are evaluated using the separation adapter.
5.3 Performance on LongEmbed MTEB
We have evaluated our model against text-embedding-3-large, bge-m3, and our previously released model suite jina-embeddings-v2 on six long document retrieval tasks from the MTEB leaderboard. The results, presented in Table 5, demonstrate that jina-embeddings-v3 with the text-matching adapter achieves the highest average performance. These findings underscore the effectiveness of the RoPE-based positional embeddings, outperforming both the fixed positional embeddings used by bge-m3 and the ALiBi-based approach employed in jina-embeddings-v2.
Model Average NarrativeQA Needle Passkey QMSum SummScreen WikiQA
jina-embeddings-v3* 70.39 33.32 84.00 100.00 39.75 92.78 72.46
jina-embeddings-v2-base-en 58.12 37.89 54.25 50.25 38.87 93.48 73.99
text-embedding-3-large 51.30 44.09 29.25 63.00 32.49 84.80 54.16
baai-bge-m3 56.56 45.76 40.25 46.00 35.54 94.09 77.73
*: text-matching adapter
Table 5: Evaluation of nDCG@10 [%] on MTEB LongEmbed Tasks.
5.4 Retrieval Failures
Hand-Selected Failure Examples [mAP in %]
Model F1 F2 F3 F4
jina-embeddings-v2-base-en 16.67 9.09 9.09 45.45
multilingual-e5-large 46.97 45.45 27.27 81.82
jina-embeddings-v3* 46.97 45.45 27.27 9.09
jina-embeddings-v3** 62.12 45.45 45.45 81.82
Synthetic Failure Examples [nDCG@10 in %]
Model F1 F2 F3 F4
jina-embeddings-v2-base-en 37.83 93.84 46.14
multilingual-e5-large 38.47 95.17 46.57
jina-embeddings-v3* 37.63 93.04 45.91
jina-embeddings-v3** 45.78 98.78 47.62
F1: Misleading Syntactic Similarities, F2: Misinterpretation of Named Entities, F3: No Understanding of Polar Questions, F4 Preference for Low Quality Documents
*: pair training, **: retrieval adapter
Table 6: Evaluation of Failure Cases
We conducted an analysis of retrieval failures typically observed when applying embedding models to retrieval tasks. This led to the identification of the four failure cases described in Section 4.3.3. To assess whether training our retrieval adapter using synthetic and preference learning datasets mitigates these failures, we performed two quantitative evaluations.
The first experiment evaluated whether failure cases in existing retrieval benchmarks, such as HotpotQA Yang et al. (2018) and NaturalQuestions Kwiatkowski et al. (2019), were resolved. These examples [^5] https://huggingface.co/datasets/jinaai/retrieval-failure-examples consist of a query, a relevant document, and a less relevant document that is often assigned a higher retrieval score. Table 6 presents the mean average precision (mAP) scores, showing that our model, after training with the retrieval adapters, handles these failure cases better or at least as effectively as our previously published jina-embeddings-v2 models and the multilingual-e5 model.
However, training on synthetic data does not seem to improve the model’s handling of failure case F2 (named entities), and failure cases F2 and F4 (low-quality documents) are handled equally well by the multilingual-e5 model. Given the small size of the evaluation sets (fewer than 10 examples for most failure cases), we conducted a second analysis using a larger, synthetically generated set of challenging examples representing typical failures. In this case, failure case F4 was excluded, as LLMs are not suited to generating low-quality content due to their training on high-quality data.
In this second experiment, our model outperforms the other three models across all tasks, as shown in the lower part of Table 6. One limitation of this evaluation approach is that the synthetic examples may be too closely aligned with the training data, potentially making these failure cases easier for the model to resolve.
5.5 Ablation Studies
5.5.1 Matryoshka Representation Learning
Table 7 presents the impact of MRL on the performance of our model. MRL is designed to improve the scalability and efficiency of text embeddings by enabling strong performance across a range of embedding dimensions (from 32 to 1024 in this case). The table is divided into two task categories: Retrieval and Semantic Textual Similarity (STS). A full evaluation can be found in Appendix A2.
In the Retrieval tasks reported in Appendix A2, our model consistently demonstrates strong performance across languages and datasets while using MRL, achieving competitive results even at lower-dimensional embeddings. For instance, in the XPQA Retrieval (French) task, the model reaches its highest score of 77.75 with a 1024-dimensional embedding, but also performs well with lower-dimensional embeddings, scoring 74.29 at 64 dimensions, representing only a 3.46% decrease. This highlights the robustness of MRL in maintaining high performance without requiring the largest embedding size.
MRL Dimension Retrieval STS
32 52.54 76.35
64 58.54 77.03
128 61.64 77.43
256 62.72 77.56
512 63.16 77.59
768 63.30 77.59
1024 63.35 77.58
Table 7: MRL ablation study on varying embedding sizes. Retrieval scores are nDCG@10 [%], STS scores are spearman correlation coefficients [%]. Full task list in Appendix A2.
5.5.2 Encoding Asymmetry for Retrieval
Task w/ Instr. w/o Instr.
1 Ad. 2 Ad. 1 Ad. 2 Ad.
SciFact 69.66 70.18 71.23 70.58
FiQA2018 46.70 47.21 47.41 47.49
TRECCOVID 73.70 75.05 60.51 70.52
NFCorpus 34.95 35.34 34.57 34.76
SCIDOCS 19.27 19.15 20.23 19.03
Touche2020 25.33 27.58 24.24 29.00
NarrativeQA 37.07 35.53 37.43 35.33
NQ 59.99 62.84 62.02 63.12
DBPedia 40.12 40.92 37.64 40.71
Average 45.20 45.98 43.92 45.62
Table 8: Evaluation of Input Type Encoding for Asymmetric Retrieval Tasks [nDCG@10 %]
Table 8 provides key insights into the impact of using one versus two adapters, as well as the influence of instructions. For most tasks, the combination of two adapters with instructions resulted in the highest performance, demonstrating the advantages of increased model capacity and explicit guidance. For instance, the highest scores for TRECCOVID and NFCorpus were achieved with two adapters and instructions, scoring 75.05 and 35.34, respectively. In contrast, the absence of instructions generally led to reduced performance, particularly when using a single adapter. This trend underscores the importance of instructions in enhancing retrieval effectiveness.
On average, the use of two adapters consistently outperformed single adapters across both instruction settings, with average scores of 45.98 and 45.62, respectively, compared to 45.20 and 43.92 for single adapters. While instructions positively contributed to performance, the increased model capacity from using two adapters had a more significant impact. These findings suggest that for optimal performance in retrieval tasks, utilizing more adapters in conjunction with instructions is generally beneficial, though task-specific factors may influence the effectiveness of these configurations.
6 Conclusion
This paper introduces jina-embeddings-v3, our latest text embedding model. By leveraging task-specific adapter tuning and failure-aware synthetic data augmentation on top of a robust backbone, jina-embeddings-v3 demonstrates competitive performance across a wide range of tasks. Extensive evaluations on both English and multilingual datasets highlight the model’s strong performance while maintaining a reasonable parameter size.
We are particularly interested in evaluating and improving the model’s performance on low-resource language and further analyzing systematic failures caused by low data availability. We plan to focus on this area going forward, further strengthening its capabilities in multilingual tasks where data availability is limited.
Appendix A Appendix
Figure 2: Scaling law of embedding models. Each dot represents an embedding model. jina-embeddings-v3 demonstrates superior performance compared to models of similar size, showing a superlinear improvement over its predecessor, jina-embeddings-v2. This graph was created by selecting 100 embedding models from the MTEB leaderboard [^7] https://huggingface.co/spaces/mteb/leaderboard, excluding those without size information, typically closed-source or proprietary models. Submissions identified as outliers or trolling were also filtered out. The average MTEB performance on English tasks is plotted against the number of model parameters. The trendline, representing all models, is highlighted, with multilingual models emphasized in orange.
Pre-Training Number of Devices Steps Batch Size Seq. Length LR*
Pre-Training (Short Text) $8$ $100,000$ $128\times 8$ 512 $1\times{}10^{-4}$
Pre-Training (Long Text) $8$ $60,000$ $8\times 8$ $8192$ $5\times{}10^{-5}$
Pair Training Number of Devices Steps Batch Size Seq. Length LR* $\tau$
Pair Training (Short Text) 8 60,000 $8\times 2048$ 192 $3\times 10^{-5}$ $0.05$
Pair Training (Long Text) 2 50,000 $2\times 512$ 1024 $2\times 10^{-5}$ $0.02$
Adapter Training Number of Devices Steps Batch Size Seq. Length LR* $\tau$
Retrieval Adapters 1 20,000 128 512 $5\times 10^{-4}$ 0.05
Text Matching Adapter 1 15,000 256 192 $1\times 10^{-4}$ 0.05
Classification Adapter 1 11,500 256 192 $5\times 10^{-4}$ 0.02
Separation Adapter 1 12,500 512 192 $5\times 10^{-5}$ 0.02
* as we use a linear learning rate scheduler with warmup, this refers to the maximum learning rate
Table A1: Hyperparameters
Task 32 64 128 256 512 768 1024
Retrieval
ArguAna (en) 45.09 50.94 52.05 53.45 54.24 54.13 54.33
QuoraRetrieval (en) 86.21 87.95 88.63 88.92 88.99 89.03 89.09
SciFact (en) 55.07 64.94 69.95 71.52 72.47 72.05 72.31
SCIDOCS (en) 14.10 16.94 18.50 19.30 19.84 19.87 19.81
TRECCOVID (en) 65.74 73.02 76.92 77.07 76.04 77.36 77.72
FiQA2018 (en) 35.31 41.02 44.51 44.64 47.29 47.36 47.35
NFCorpus (en) 25.92 31.62 34.7 35.79 36.21 36.55 36.62
AlloprofRetrieval (fr) 40.19 46.91 51.22 53.35 53.98 54.39 54.28
BSARDRetrieval (fr) 15.79 19.85 23.55 24.48 24.92 25.15 24.76
XPQARetrieval (fr) 67.81 74.29 75.87 77.36 77.46 77.31 77.59
CmedqaRetrieval (zh) 28.67 32.20 34.16 35.32 35.85 35.82 35.93
DuRetrieval (zh) 73.65 79.36 81.66 82.83 83.11 83.15 83.13
EcomRetrieval (zh) 38.40 49.65 57.59 59.75 60.86 60.73 60.89
MedicalRetrieval (zh) 44.24 51.76 54.74 56.27 56.36 56.57 56.64
GermanDPR (de) 75.61 79.59 81.82 82.23 82.45 82.34 82.42
GermanQuAD-Retrieval (de) 90.00 93.13 94.78 95.29 95.47 95.48 95.46
MLQARetrieval (es) 51.95 59.66 64.66 67.49 68.32 68.61 68.63
SpanishPassageRetrievalS2S (es) 62.18 66.71 70.57 70.14 69.40 70.05 69.83
XPQARetrieval (es) 61.47 68.25 70.26 71.33 71.71 71.91 72.01
RiaNewsRetrieval (ru) 62.84 72.26 76.66 78.58 78.99 79.11 79.21
RuBQRetrieval (ru) 63.05 69.28 71.55 72.11 72.49 72.39 72.31
STS
BIOSSES (en) 87.75 88.00 88.04 88.04 88.45 88.66 88.69
SICK-R (en) 88.77 89.47 89.65 89.70 89.69 89.66 89.62
STS12 (en) 79.41 81.32 82.23 82.41 82.40 82.43 82.44
STS15 (en) 85.84 87.94 89.12 89.24 89.22 89.28 89.31
STS22 (en) 67.52 67.29 67.51 67.37 67.44 67.51 67.28
STSBenchmarkMultilingualSTS (de) 85.37 86.76 87.62 87.83 87.83 87.82 87.80
STS22 (de) 61.43 61.01 61.45 61.54 61.66 61.74 61.61
STSBenchmarkMultilingualSTS (es) 85.71 87.09 87.78 87.92 88.02 88.00 87.96
STS22 (es) 69.59 71.22 72.00 72.04 72.00 72.17 72.21
STSBenchmarkMultilingualSTS (fr) 85.80 87.00 87.35 87.41 87.42 87.43 87.44
STS22 (fr) 83.84 83.29 82.78 83.40 83.50 83.16 83.29
SICKFr (fr) 83.19 83.83 83.83 83.94 83.95 83.92 83.95
STSBenchmarkMultilingualSTS (zh) 84.73 85.33 85.66 85.69 85.72 85.69 85.64
STS22 (zh) 66.18 66.15 66.26 66.22 66.15 66.19 66.23
AFQMC (zh) 42.09 42.93 43.77 43.76 43.54 43.48 43.47
STSBenchmarkMultilingualSTS (ru) 84.81 85.69 86.01 86.30 86.34 86.30 86.30
STS22 (ru) 66.36 65.20 65.11 65.22 65.08 65.12 65.09
STSBenchmarkMultilingualSTS (pl) 82.55 83.99 85.18 85.52 85.58 85.50 85.48
STS22 (pl) 45.04 45.30 45.01 45.14 45.17 45.10 45.14
CDSC-R (pl) 91.03 91.87 92.33 92.48 92.59 92.62 92.63
Table A2: MLR ablation study on retrieval and STS tasks
jina-embeddings-v3 text-embedding-3 multilingual-e5 cohere-embed
-large -large-instruct -multilingual-v3
Retrieval Average [nDCG@10 %] 53.87 55.44 52.47 53.84
ArguAna (en) 54.33 58.05$\star$ 58.38 55.11
CQADupstackRetrieval (en) 42.36 47.54$\star$ 42.71 40.64
ClimateFEVER (en) 42.36 30.27 29.86 29.96
DBPedia (en) 41.00 44.76 38.36 41.00
FEVER (en) 89.06 87.94 77.99 88.53
FiQA2018 (en) 47.35 55.00 47.71 44.10
HotpotQA (en) 64.67 71.58 69.32 70.61
MSMARCO (en) 40.82 40.24 40.43 43.45
NFCorpus (en) 36.62 42.07 35.53 36.42
NQ (en) 64.23 61.27 57.75 63.41
QuoraRetrieval (en) 89.09 89.05$\star$ 89.15 88.92
SCIDOCS (en) 19.81 23.11 18.72 19.34
SciFact (en) 72.31 77.77 71.85 70.05
Touche2020 (en) 26.30 23.35 27.25 32.70
TRECCOVID (en) 77.72 79.56 82.00 83.37
Clustering Average [v–measure] 45.27 49.00 47.10 46.60
ArxivClusteringP2P (en) 46.66 49.01 46.40 48.16
ArxivClusteringS2S (en) 39.27 44.45 40.49 40.79
BiorxivClusteringP2P (en) 38.91 38.03 40.94 40.50
BiorxivClusteringS2S (en) 34.42 36.53 36.28 36.91
MedrxivClusteringP2P (en) 34.80 32.70 36.93 36.18
MedrxivClusteringS2S (en) 32.42 31.27 35.54 33.44
RedditClustering (en) 55.40 67.84 56.60 58.11
RedditClusteringP2P (en) 62.87 67.96 64.27 65.02
StackExchangeClustering (en) 65.66 76.26 66.85 68.12
StackExchangeClusteringP2P (en) 35.08 36.88 42.46 35.22
TwentyNewsgroupsClustering (en) 52.49 58.14 51.33 50.14
Reranking Average [map] 58.13 59.16 58.58 57.86
AskUbuntuDupQuestions 65.04 65.03 63.89 62.13
MindSmallReranking 30.83 29.86 33.09 32.59
SciDocsRR 84.88 86.66 85.87 84.31
StackOverflowDupQuestions 51.77 55.08 51.45 52.40
Classification Average [acc.] 82.58 75.45 77.56 76.01
AmazonCounterfactualClassification 89.49 78.83 76.24 77.85
AmazonPolarityClassification 95.38 92.85 96.29 95.60
AmazonReviewsClassification 49.77 48.70 56.72 49.79
Banking77Classification (en) 84.08 85.69 85.73 86.09
EmotionClassification (en) 73.30 51.58 51.51 48.15
ImdbClassification (en) 91.90 87.67 94.60 93.97
MTOPDomainClassification (en) 97.49 95.36 93.93 94.92
MTOPIntentClassification (en) 84.53 75.05 82.46 78.89
MassiveIntentClassification (en) 77.60 74.64 77.06 74.51
MassiveScenarioClassification (en) 84.71 79.79 80.47 79.00
ToxicConversationsClassification (en) 91.28 72.92 71.06 71.20
TweetSentimentExtractionClassification (en) 71.39 62.22 64.62 62.18
Pair–Classification Average [AveP] 84.01 86.19 86.15
SprintDuplicateQuestions (en) 96.98 91.18 96.79
TwitterSemEval2015 (en) 70.91 80.27 75.16
TwitterURLCorpus (en) 84.13 87.12 86.49
STS Average [Spearman] 85.80 84.78 83.15
BIOSSES 88.69 86.96 85.01
SICK–R 89.62 81.73 82.18
STS12 82.44 82.57 77.62
STS13 89.49 87.15 85.16
STS14 84.94 84.97 80.02
STS15 89.31 91.05 88.92
STS16 86.84 87.31 86.92
STS17 (en–en) 89.97 90.03 90.09
STS22 (en) 67.27 67.63 66.81
STSBenchmark 89.44 88.38 88.79
Summarization Average [Spearman] 29.7 30.39 30.99
SummEval 29.70 30.39 30.99
$\star$: ArguAna, CQADupstackRetrieval and QuoraRetrieval are evaluated using the text matching adapter for jina-embeddings-v3
Table A3: Performance of embedding models on English MTEB tasks.
je-v2* je-v3 m-e5-large te-3-large Cohere-embed m-e5-instr
Retrieval Average [nDCG@10] 58.24 57.98 52.37 58.38
MIRACL (all) 61.9 66.5 54.9 52.8 65.7
Chinese Average 69.40 68.60 63.66 64.94
CmedqaRetrieval 39.15 35.93 28.67 34.11
CovidRetrieval 81.22 78.92 75.51 75.76
DuRetrieval 84.57 83.13 85.32 85.14
EcomRetrieval 63.95 60.89 54.75 53.94
MedicalRetrieval 57.12 56.64 51.44 56.56
MMarcoRetrieval 77.96 79.69 79.2 78.82
T2Retrieval 80.59 83.16 76.11 82.92
VideoRetrieval 70.62 70.43 58.25 52.28
German Average 38.73 42.32 35.63 38.19
GerDaLIR 17.20 16.23 6.53 9.33
GermanDPR 79.50 82.42 82.89 80.87
XMarket 19.50 28.32 17.46 24.37
Spanish Average 52.12 47.75 44.55 47.86
MintakaRetrieval 28.30 26.91 33.23 34.50
XMarket 19.70 26.59 13.48 25.00
SpanishPassageRetrievalS2S 81.07 69.83 72.32 71.49
SpanishPassageRetrievalS2P 66.15 43.4 41.96 43.17
XPQARetrieval 65.40 72.00 61.77 65.12
French Average 53.48 42.17 40.42 62.46
AlloprofRetrieval 54.28 38.15 38.36 52.07
BSARDRetrieval 24.76 0.27 0.14 66.21
MintakaRetrieval (fr) 26.91 25.20 25.44 33.49
SyntecRetrieval 83.85 81.07 79.27 87.78
XPQARetrieval (fr) 77.58 66.15 58.87 72.73
Russian Average 75.76 77.39 78.57
RiaNewsRetrieval (rus–Cyrl) 79.21 80.67 83.25
RuBQRetrieval (rus–Cyrl) 72.30 74.11 73.88
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   te-3-large: text-embedding-3-large
Cohere-embed: Cohere-embed-multilingual-v3   m-e5-instr: multilingual-e5-large-instruct
Table A4: Performance of bi-/multilingual models on multilingual retrieval tasks.
je-v2* je-v3 m-e5-large m-e5-instr
STS Average [Spearman %] 66.60 69.83 64.65 68.77
Chinese Average 59.38 54.18 48.29 53.63
AFQMC 50.59 43.47 33.02 37.54
ATEC 51.28 49.06 39.81 43.26
BQ 66.07 63.2 46.44 48.80
LCQMC 75.74 75.95 75.95 76.05
PAWSX 41.48 15.56 14.63 15.02
QBQTC 38.11 36.43 29.77
STS22 69.25 66.23 65.64 73.10
STSB 82.55 83.54 81.08 81.67
German Average 78.47 78.97 74.83 77.22
STSBenchmarkMultilingualSTS 88.45 87.8 84.27 85.37
GermanSTSBenchmark 88.32 87.5 83.64 84.84
STS22 58.63 61.61 56.59 61.45
Spanish Average 77.72 80.09 74.21 79.93
STSBenchmarkMultilingualSTS 86.38 87.96 83.81 86.15
STS22 69.06 72.21 64.60 73.70
French Average 84.89 79.37 82.62
STSBenchmarkMultilingualSTS 87.44 76.79 84.93
STS22 83.29 82.53 82.73
SICKFr 83.95 78.78 80.20
Polish Average 72.76 70.45 72.15
CDSC–R 92.63 91.00 92.35
SICK–R–PL 80.53 75.08 77.61
STS22 45.14 34.66 46.49
Russian Average 81.5 74.48 79.67
RUParaPhraserSTS 76.77 71.82 75.37
RuSTSBenchmarkSTS 86.22 83.15 83.97
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   m-e5-instr: multilingual-e5-large-instruct
Table A5: Performance of bi-/multilingual models on STS tasks.
je-v2* je-v3 m-e5-large m-e5-instr
Pair Classification Average [AveP] 82.95 76.91 76.95 77.79
Chinese Average 82.95 72.76 69.89 66.54
Cmnli 85.27 79.10 78.18 71.41
Ocnli 80.62 66.42 61.6 61.67
French Average 76.57 76.19 77.23
OpusparcusPC 93.74 93.89 94.72
PawsXPairClassification 59.4 58.5 59.73
Polish Average 83.70 85.5 87.16
CDSC–E 73.02 74.47 76.18
PpcPC 91.43 92.18 93.45
PSC 99.19 99.39 99.31
SICK–E–PL 71.15 75.96 79.68
Russian Average 58.77 58.4 63.92
TERRa 58.77 58.4 63.92
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   m-e5-instr: multilingual-e5-large-instruct
Table A6: Performance of bi-/multilingual models on pair-classification tasks.
je-v2* je-v3 m-e5-large m-e5-instr
Classification Average [acc.] 65.69 71.46 65.22 67.45
Chinese Average 64.94 69.07 67.34 67.85
AmazonReviews 34.94 44.77 38.83 45.11
IFlyTek 47.36 41.68 45.47 44.06
JDReview 79.57 83.51 80.99 80.21
MassiveIntent 68.2 73.22 71.12 67.85
MassiveScenario 71.93 80.82 76.83 72.43
MultilingualSentiment 63.29 73.21 68.58 72.44
OnlineShopping 87.00 91.88 90.81 91.83
TNews 47.65 45.97 48.38 49.84
Waimai 84.54 86.59 85.02 86.85
German Average 65.66 79.03 67.77 69.55
AmazonCounterfactual 68.92 89.61 68.66 65.63
AmazonReviews 37.72 48.07 46.50 54.54
MassiveIntent 63.89 74.84 63.82 65.90
MassiveScenario 71.25 83.97 71.25 72.69
MTOPDomain 88.37 96.89 90.44 90.00
MTOPIntent 63.83 80.79 65.97 68.55
Spanish Average 67.10 77.59 65.33 66.89
AmazonReviews 38.68 48.10 44.35 49.88
MassiveIntent 66.93 75.32 64.01 65.57
MassiveScenario 71.23 82.58 69.07 70.01
MTOPDomain 89.89 97.17 88.34 89.12
MTOPIntent 68.76 84.77 61.90 69.86
French Average 76.70 68.39 69.32
AmazonReviews 47.3 41.91 49.78
MasakhaNEWS 74.99 79.38 78.93
MassiveIntent 76.33 69.34 66.88
MassiveScenario 83.09 73.87 71.16
MTOPDomain 96.30 86.41 85.89
MTOPIntent 82.21 59.43 63.29
Polish Average 70.81 63.82 65.99
AllegroReviews 49.51 41.14 45.32
CBD 69.99 69.90 74.25
MassiveIntent 75.37 65.07 67.45
MassiveScenario 82.1 69.82 71.44
PAC 68.4 70.37 65.69
PolEmo2.0–IN 83.75 77.06 80.99
PolEmo2.0–OUT 66.53 53.38 56.84
Russian Average 59.84 58.92 65.38
eoreviewClassification 48.01 49.69 55.92
HeadlineClassification 75.08 77.19 86.18
Inappropriateness 61.05 61.60 65.61
RuSciBenchGRNTI 59.19 58.20 65.08
RuSciBenchOECD 45.56 43.91 50.18
Kinopoisk 62.39 56.59 66.12
RuReviews 67.58 65.28 68.54
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   te-3-large: text-embedding-3-large
Cohere-embed: Cohere-embed-multilingual-v3   m-e5-instr: multilingual-e5-large-instruct
Table A7: Performance of bi-/multilingual models on classification tasks.
je-v2* je-v3 m-e5-large te-3-large m-e5-instr
Clustering Average [v–measure] 39.36 46.71 42.12 46.65 52.12
Chinese Average 46.47 50.22 48.23 48.75 52.72
CLSClusteringS2S 38.4 38.79 38.59 38.82 40.65
CLSClusteringP2P 39.97 39.96 40.68 39.64 42.28
ThuNewsClusteringS2S 53.42 58.98 55.59 55.40 60.81
ThuNewsClusteringP2P 54.08 63.13 58.05 61.15 67.12
German Average 29.91 36.54 34.60 37.26 39.73
TenKGnadClusteringS2S 25.01 39.76 33.9 37.25 41.42
TenKGnadClusteringP2P 43.07 43.76 45.8 45.03 52.76
BlurbsClusteringS2S 16.67 21.00 16.51 21.98 20.99
BlurbsClusteringP2P 34.89 41.63 42.19 44.78 43.78
Spanish Average 44.05 45.09 41.52 48.59 46.50
SpanishNewsClusteringP2P 48.31 45.04 41.66 52.96 45.98
FloresClusteringS2S 39.79 45.13 41.38 44.21 47.01
French Average 44.95 38.63 45.79 55.80
HALClusteringS2S 27.48 22.44 27.87 28.25
AlloProfClusteringS2S 46.86 32.26 53.56 60.03
AlloProfClusteringP2P 63.88 62.99 62.72 68.77
MLSUMClusteringP2P 44.69 44.04 45.01 47.25
MLSUMClusteringS2S 44.88 37.65 38.41 46.18
MasakhaNEWSClusteringP2P 42.36 40.49 53.23 70.48
MasakhaNEWSClusteringS2S 44.48 30.56 39.71 69.65
Russian Average 60.80 52.55 57.10 63.04
GeoreviewClusteringP2P 73.62 60.51 71.59 74.29
RuSciBenchGRNTIClusteringP2P 57.99 52.03 53.69 61.74
RuSciBenchOECDClusteringP2P 50.79 45.11 46.01 53.09
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   te-3-large: text-embedding-3-large
m-e5-instr: multilingual-e5-large-instruct
Table A8: Performance of bi-/multilingual models on clustering tasks.
je-v2* je-v3 m-e5-large m-e5-instr
Reranking Average 66.57 63.98 63.40 69.02
Chinese Average 66.57 60.64 56.00 60.68
CMedQAv1 83.64 77.96 68.25 75.94
CMedQAv2 83.74 78.22 68.56 76.10
MMarcoReranking 31.54 21.05 21.34 23.59
T2Reranking 67.37 65.31 65.83 67.10
French Average 69.86 72.14 82.30
AlloprofReranking 66.39 57.37 74.65
SyntecReranking 73.32 86.9 89.95
Russian Average 65.57 75.58 75.85
RuBQReranking 65.57 75.58 75.85
je-v2*: jina-embeddings-v2-base-(zh/de/es) bilingual model suite   je-v3: jina-embeddings-v3
m-e5-large: multilingual-e5-large   m-e5-instr: multilingual-e5-large-instruct
Table A9: Performance of bi-/multilingual models on reranking tasks.